OpenStudy (anonymous):

identify the 31st term of an arithmetic sequence where a1= 26 and a22 = -226

OpenStudy (anonymous):

a22 = a1 + 21d -226 = 26 + 21d 21d = -252 d = -12 31st term = 26 + (-12)30 = 26 - 360 = - 334

OpenStudy (anonymous):

thankk you

OpenStudy (campbell_st):

find the common difference using the 22 term use the term formula \[T _{n} = a + (n -1)d\] -226 = 26 + (22-1)d -226 = 26 + 21d 21d = -252 d = -12 then for term 31 \[T _{31} = 26 + (31 -1)\times -12\] \[T _{31} = 26 - 360\] \[T _{31} = - 334\]

OpenStudy (anonymous):

thats ok - i hope you can now be more confident in solving these practice some

OpenStudy (anonymous):

im trying, its just hard teaching myself, also im not good with word problems